# Winding quotients for virtual period maps of rank 1

### v3

@@ 6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
whenever $\Im(u+m\tau)$ is non-zero for all integers $m$, where the sign $\pm$ means ++ if $|\exp(2\pi\sqrt{-1}n(u+m\tau))|>1$ and - if $|\exp(2\pi\sqrt{-1}n(u+m\tau))|<1$. That is, the formula holds on the open set $\mathbb{C}_{u}\times\mathbb{H}_{\tau}\setminus\cup_{m\in\mathbb{Z}}\{\Im(u+m\tau)=0\}$. Clearly, it projects onto the defining domain $\mathbb{H}_{\tau}$ of $E_{2}(\tau)$. ∎
@@
Yosuke Ohyama has informed author that the function $\mathcal{W}(w,q)$ can be expressed by a bilateral basic hypergeometric series ([G-R] chap.5.) as follows;
@@ 7. Solution to the new inversion problem
Acknowledgement: The author express his gratitude to Hiroki Aoki who informed him the Borcherds work and helped him in solving the new inversion problem, and to Akishi Ikeda who explained him the roles of the pull-back of the function $\mathcal{W}$ to the winding covering space in mathematical physics and informed him its related literatures (see Remark 1.1). He also express his gratitude to Masahiko Yoshinaga for discussions during the preparation of the present note.

### v4

@@ 6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
whenever $\Im(u+m\tau)$ is non-zero for all integers $m$, where the sign $\pm$ means ++ if $|\exp(2\pi\sqrt{-1}n(u+m\tau))|<1$ and - if $|\exp(2\pi\sqrt{-1}n(u+m\tau))|>1$.
That is, the pull-back relation holds on the open set $\mathbb{C}_{u}\times\mathbb{H}_{\tau}\setminus\cup_{m\in\mathbb{Z}}\{\Im(u+m\tau)=0\}$. Clearly, this set projects surjectively onto $\mathbb{H}_{\tau}$. ∎
@@
Yosuke Ohyama has informed author that the function $\mathcal{W}(w,q)$ can be expressed by a bilateral basic hypergeometric series ([GR] chap.5.) as follows;
@@ 7. Solution to the new inversion problem
Acknowledgements. The author express his gratitude to Hiroki Aoki who informed him the Borcherds work (see the proof of Assertion B) and helped him in solving the new inversion problem, and to Akishi Ikeda who explained him the roles of the pull-back of the function $\mathcal{W}$ to the winding covering space in mathematical physics and informed him its related literatures (see Remark 1.1). He also express his gratitude to Masahiko Yoshinaga for discussions during the preparation of the present note.